# Superposition in Graph Neural Networks

### v1

@@ 5 Pooling-Induced Axis Alignment of Graph Representations
Metric and preprocessing. We measure axis preference with the Alignment Index (AI) computed on any set of feature directions $\{c_{\ell}\in\mathbb{R}^{d}\}_{\ell=1}^{k}$ (node-level probe normals or graph-level class features):
@@
Thus, $\mathrm{AI}\approx 1/\sqrt{d}$ for random orientations and $\mathrm{AI}\to 1$ when each feature is concentrated on a coordinate axis. We always use row-unit directions; for centroids we COM-center across classes before normalizing (to remove the shared offset), while probes use the raw fitted normals.
Global and local pooling conditions. For the graph readout we use the signed power-mean family that interpolates mean ($p=1$) and max ($p\to\infty$):
@@
To isolate local effects, we also replace the last message-passing aggregator by the same power-mean operator while holding the global readout at mean pooling.
Alignment vs. pooling and regime. Across both a bottleneck ($d=10$) and a wide ($d=22$) regime, AI increases as p>1p\!>\!1 (Fig. 4). With global pooling the effect is clearest at the graph level (the nonlinearity acts after node mixing). With a local power-mean in the last layer, AI increases at both node and graph levels because the winner-take-most bias precedes the final ReLU and readout. Occasionally AI dips for p>2p\!>\!2 in the bottleneck regime; this coincides with less stable training and noisier features. The trends persist when conditioning on high-accuracy runs. GATv2 tends to exhibit higher node-level AI even for modest $p_{\text{local}}$ (attention already induces axis-selective flows), whereas GCN’s degree-normalized linear aggregation produces smoother trends.
From alignment to constrained superposition. Under mean pooling, strong bottlenecks yield large $\mathrm{SI}$ due to shared directions, especially at the node level where separation is not directly optimized by the loss. Increasing $p_{\text{local}}$ reliably reduces $\mathrm{SI}_{\text{node}}$ (Fig. 4), effectively capping features-per-axis. Raising $p_{\text{global}}$ also lowers $\mathrm{SI}_{\text{node}}$ in the wide regime (consistent with §4.3); in the bottleneck regime the dimensional constraint dominates and this effect largely vanishes. $\mathrm{SI}_{\text{graph}}$ stays near-flat: in the wide regime it is already close to optimal under mean pooling (§4.2), and in the bottleneck regime additional alignment cannot overcome the shortage of dimensions.
@@
Figure 4: Alignment and superposition indices versus pooling parameter $p$ across architectures.
Noise as the driver of axis alignment. Two complementary arguments explain why increasing $p>1$ breaks rotational symmetry in favor of axis alignment. (i) Learning-dynamics view. Backpropagating through power-mean pooling yields
$\frac{\partial h_{d}}{\partial x_{id}}\;=\;\frac{1}{N^{1/p}\,|s_{d}|^{1/p-1}}\operatorname{sgn}(s_{d})\,|x_{id}|^{p-1}\operatorname{sgn}(x_{id}),$
so large-magnitude coordinates get disproportionally large updates. With $p=1$ the Jacobian is rotation-invariant; with $p>1$ the extra factor $|x_{id}|^{p-1}$ biases gradient flow toward coordinates with larger magnitude, steadily aligning features to axes. In the absence of noise all orientations become equivalent once the model has recognized the features; in realistic graphs ubiquitous irrelevant features inject noise and trigger the symmetry breaking. (ii) Geometric view. Under equal-energy corruption, axis-aligned vectors lose less information than randomly oriented ones; in high dimensions the components of a random unit vector scale like $1/\sqrt{d}$ and are easily swamped by noise spikes. Figure 5 (from our diagnostic) visualizes this effect under max pooling. Empirically, purely noisy inputs confirm the variance story: max pooling inflates the standard deviation of pooled embeddings by an order of magnitude relative to mean pooling.
Practical takeaways. For more axis-aligned graph features, prefer global pooling with $p_{\text{global}}>1$. To make node channels less polysemantic, consider a power-mean local aggregator in the final layer.
Figure 5: Axes-aligned embedding vectors lose less information than arbitrary-angled vectors under max pooling when both are corrupted by equal-energy noise.
@@ 6 Rank collapse and metastable superposition
If the span of $H$ is low-dimensional, superposition is inevitable: by pigeonhole, $k_{a}>\mathrm{rank}(H)$ forces multiple features to share axes. We formalize this with a numerical rank
@@
On PAIRWISE we observe two characteristic dynamics (Fig. 6). In some seeds $r_{\tau}(H)$ rises to the layer width (d=16d\!=\!16), with $\mathrm{EffRank}(H)$ tracking from below; SI/WNO on centroids drop accordingly. In other seeds $r_{\tau}(H)$ remains strictly below $d$, sometimes far below $k_{a}$: SI/WNO stay high and the smallest singular values show brief activation attempts (spikes) that recede, indicating the optimizer returns to a low-rank basin. Interestingly, across models that exhibit overfitting (such that the right model in Fig. 6) we observe sharp multi-phase transitions in SI and WNO throughout training. A possible (non-rigorous) explanation is provided in Appendix D.
@@
Why does the low-rank basin persist? Two effects make it metastable: (i) Global channel gating by ReLU. If a channel is negative across the dataset just before the last nonlinearity, ReLU zeros it everywhere, so the corresponding column of $H$ vanishes and algebraic rank drops exactly. Replacing the final ReLU with LeakyReLU substantially reduces such dead columns (App. E.6). (ii) Obtuseness pressure under BCE. Before a separating hyperplane forms, last-layer directions $\{\mathbf{v}_{i}\}$ tend to align with their class weights $\{\mathbf{w}_{i}\}$ and become mutually obtuse. Moving one feature into a fresh orthogonal dimension rotates it towards $90^{\circ}$ against the others and increases the BCE loss. After the hyperplane forms, the cross-class margins satisfy $\mathbf{w}_{j}^{\!\top}\mathbf{v}_{i}<0$ ($i\neq j$), and perturbing $\mathbf{v}_{k}$ into a new dimension makes those dots less negative, again increasing loss (Appendix E). Hence gradients oppose escapes until the accuracy gain from a new dimension outweighs the obtuseness penalty, explaining the singular jump events.
When is collapse likely? We see more collapse (i) with narrower last layers, (ii) in GIN (post-aggregation ReLUs) than GCN (no ReLU after the final aggregation), and (iii) as the number of task features grows. This agrees with the regime analysis in Appendix E: when $k_{a}\!\leq\!d+1$, mutually obtuse low-rank configurations exist and can act as local minima; as $k_{a}$ approaches or exceeds $2d$, such configurations become geometrically impossible and $r_{\tau}(H)$ tends to reach $d$.
Connection to oversmoothing. Classic oversmoothing is within-graph: node states become indistinguishable with depth. Here we see a cross-graph analogue: pooled graph embeddings concentrate in a thin global subspace. Accurate models can then separate classes inside that span by sign/magnitude, which explains why SI/WNO can be high at the node level yet acceptable in the graph space.
Discussion. The width-sweeps in §4.2 showed SI/WNO can be high even when $d\!\geq\!k$; here we identify a complementary mechanism: training can settle into thin-span solutions, so overlap is forced (because $k_{a}>r_{\tau}(H)$) rather than just suboptimal packing within a rich span. This also clarifies why node-level superposition can persist while graph-level centroids look well-packed: the readout operates inside a low-dimensional subspace, separating classes by sign/magnitude without increasing the span of $H$.
@@ 7 Conclusion
In this work, we provided, to the best of our knowledge, the first systematic study of superposition in GNNs. Concretely, we introduced a representation-centric framework that quantifies feature sharing in GNNs via class-conditional centroids and linear-probe directions. Across architectures, width induces a three-phase trajectory in overlap; topology shapes node-level geometry that pooling re-mixes toward the task; and sharper pooling drives axis alignment and reduces lever sharing. We also observed metastable low-rank graph embeddings in shallow models. Practically, modest increases in width, LeakyReLU in the final MLP, and sharper but stable pooling improve interpretability without sacrificing accuracy. A key next step is to link superposition dynamics to over-smoothing and over-squashing, integrating our geometric view with spectral, dynamical, and topological analyses.
Appendix A Generalized Mean Pooling

### v2

@@ 5 Pooling-Induced Axis Alignment of Graph Representations
Metric and preprocessing. We measure axis preference with the Alignment Index (AI) computed on any set of feature directions $\{c_{\ell}\in\mathbb{R}^{d}\}_{\ell=1}^{k}$ (node‑level probe normals or graph‑level class features):
@@
Thus, $\mathrm{AI}\approx 1/\sqrt{d}$ for random orientations and $\mathrm{AI}\to 1$ when each feature is concentrated on a coordinate axis. We always use row‑unit directions; for centroids we COM‑center across classes before normalizing (to remove the shared offset), while probes use the raw fitted normals.
Global and local pooling conditions. For the graph readout we use the signed power‑mean family (details in Appendix C) that interpolates mean ($p=1$) and max ($p\to\infty$):
@@
To isolate local effects, we also replace the last message‑passing aggregator by the same power‑mean operator while holding the global readout at mean pooling.
Alignment vs. pooling and regime. On Pairwise ($k=16$) across both a bottleneck ($d=10$) and a wide ($d=22$) regime, AI increases as p>1p\!>\!1 (Fig. 4). With global pooling the effect is clearest at the graph level (the nonlinearity acts after node mixing). With a local power‑mean in the last layer, AI increases at both node and graph levels because the winner‑take‑most bias precedes the final ReLU and readout. Occasionally AI dips for p>2p\!>\!2 in the bottleneck regime; this coincides with less stable training and noisier features. The trends persist when conditioning on high‑accuracy runs. GATv2 tends to exhibit higher node‑level AI even for modest $p_{\text{local}}$ (attention already induces axis‑selective flows), whereas GCN’s degree‑normalized linear aggregation produces smoother trends.
From alignment to constrained superposition. Under mean pooling, strong bottlenecks yield large $\mathrm{SI}$ due to shared directions, especially at the node level where separation is not directly optimized by the loss. Increasing $p_{\text{local}}$ reliably reduces $\mathrm{SI}_{\text{node}}$ (Fig. 4), effectively capping features‑per‑axis. Raising $p_{\text{global}}$ also lowers $\mathrm{SI}_{\text{node}}$ in the wide regime (consistent with §4.3); in the bottleneck regime the dimensional constraint dominates and this effect largely vanishes. $\mathrm{SI}_{\text{graph}}$ stays near‑flat: in the wide regime it is already close to optimal under mean pooling (§4.2), and in the bottleneck regime additional alignment cannot overcome the shortage of dimensions.
@@
Figure 4: Alignment versus pooling parameter $p$ across architectures on Pairwise ($k=16$). Shaded regions show uncertainty as mean $\pm 1.96\sigma/\sqrt{R}$ where $R=16$ is the number of seeds.
Noise as the driver of axis alignment. Two complementary arguments explain why increasing $p>1$ breaks rotational symmetry in favor of axis alignment. (i) Learning-dynamics view. Backpropagating through generalized mean pooling (App. C) yields, for each pooled feature $h_{d}$,
$\frac{\partial h_{d}}{\partial x_{id}}\;=\;\frac{1}{N}\bigl(|\bar{s}_{d}|+\varepsilon\bigr)^{\frac{1}{p}-1}\bigl(|x_{id}|+\varepsilon\bigr)^{p-1},$
(a derivation is given in Appendix D) where $\bar{s}_{d}$ is the mean of the transformed features and $\varepsilon>0$ is the stabilizer. For fixed $d$ and $p$, the prefactor $\frac{1}{N}\bigl(|\bar{s}_{d}|+\varepsilon\bigr)^{\frac{1}{p}-1}$ is shared across nodes, so the relative gradient magnitudes are controlled by $(|x_{id}|+\varepsilon)^{p-1}$: with $p=1$ we recover $\partial h_{d}/\partial x_{id}=1/N$ (a rotation-invariant Jacobian), whereas for $p>1$ coordinates with larger $|x_{id}|$ receive disproportionately larger updates, steadily aligning features to coordinate axes. In the absence of noise all orientations become equivalent once the model has recognized the features; in realistic graphs ubiquitous irrelevant features inject noise and trigger the symmetry breaking.
(ii) Geometric view. Under equal-energy corruption, axis-aligned vectors lose less information than randomly oriented ones; in high dimensions the components of a random unit vector scale like $1/\sqrt{d}$ and are easily swamped by noise spikes. Figure 5 visualizes this effect under max pooling. We numerically simulate the equal-energy corruption under max pooling: for each $(n,\sigma)$ on a grid (dimensions $n=2\ldots 20$, noise levels $\sigma\in[0.001,0.35]$) we sample $100$ random vectors $v\in\mathbb{R}^{n}$ (either uniform on the unit sphere or one-hot), draw $20$ i.i.d. Gaussian noise candidates per coordinate with variance $\sigma^{2}$, and replace each coordinate by the candidate with largest absolute value whenever its magnitude exceeds $|v_{i}|$.
Practical takeaways. For more axis‑aligned graph features, prefer global pooling with $p_{\text{global}}>1$. To make node channels less polysemantic, consider a power‑mean local aggregator in the final layer.
Figure 5: Axes‑aligned embedding vectors lose less information than arbitrary‑angled vectors under max pooling when both are corrupted by equal‑energy noise.
@@ 6 Rank collapse and metastable superposition
If the span of $H$ is low‑dimensional, superposition is inevitable: by pigeonhole, $k_{a}>\mathrm{rank}(H)$ forces multiple features to share axes. We formalize this with a numerical rank
@@
On PAIRWISE we observe two characteristic dynamics (Fig. 6). In some seeds $r_{\tau}(H)$ rises to the layer width (d=16d\!=\!16), with $\mathrm{EffRank}(H)$ tracking from below; SI/WNO on centroids drop accordingly. In other seeds $r_{\tau}(H)$ remains strictly below $d$, sometimes far below $k_{a}$: SI/WNO stay high and the smallest singular values show brief activation attempts (spikes) that recede, indicating the optimizer returns to a low‑rank basin. Interestingly, across models that exhibit overfitting (such that the right model in Fig. 6) we observe sharp multi-phase transitions in SI and WNO throughout training. A possible (non-rigorous) explanation is provided in Appendix E.
@@
Why does the low‑rank basin persist? Two effects make it metastable: (i) Global channel gating by ReLU. If a channel is negative across the dataset just before the last nonlinearity, ReLU zeros it everywhere, so the corresponding column of $H$ vanishes and algebraic rank drops exactly. Replacing the final ReLU with LeakyReLU substantially reduces such dead columns (App. F.5). (ii) Obtuseness pressure under BCE. Before a separating hyperplane forms, last‑layer directions $\{\mathbf{v}_{i}\}$ tend to align with their class weights $\{\mathbf{w}_{i}\}$ and become mutually obtuse. Moving one feature into a fresh orthogonal dimension rotates it towards $90^{\circ}$ against the others and increases the BCE loss. After the hyperplane forms, the cross‑class margins satisfy $\mathbf{w}_{j}^{\!\top}\mathbf{v}_{i}<0$ ($i\neq j$), and perturbing $\mathbf{v}_{k}$ into a new dimension makes those dots less negative, again increasing loss (Appendix F). Hence gradients oppose escapes until the accuracy gain from a new dimension outweighs the obtuseness penalty, explaining the singular jump events.
When is collapse likely?
We see more collapse (i) with narrower last layers, (ii) in GIN (post‑aggregation ReLUs) than GCN (no ReLU after the final aggregation), and (iii) as the number of task features grows. This agrees with the regime analysis in Appendix F: when $k_{a}\!\leq\!d+1$, mutually obtuse low‑rank configurations exist and can act as local minima; as $k_{a}$ approaches or exceeds $2d$, such configurations become geometrically impossible and $r_{\tau}(H)$ tends to reach $d$.
Connection to oversmoothing. Classic oversmoothing is within‑graph: node states become indistinguishable with depth. Here we see a cross‑graph analogue: pooled graph embeddings concentrate in a thin global subspace. Accurate models can then separate classes inside that span by sign/magnitude, which explains why SI/WNO can be high at the node level yet acceptable in the graph space.
Discussion. The width‑sweeps in §4.2 showed SI/WNO can be high even when $d\!\geq\!k$; here we identify a complementary mechanism: training can settle into thin‑span solutions, so overlap is forced (because $k_{a}>r_{\tau}(H)$) rather than just suboptimal packing within a rich span. This also clarifies why node‑level superposition can persist while graph‑level centroids look well‑packed: the readout operates inside a low‑dimensional subspace, separating classes by sign/magnitude without increasing the span of $H$.
@@ 7 Conclusion
In this work, we provided, to the best of our knowledge, the first systematic study of superposition in GNNs. Concretely, we introduced a representation‑centric framework that quantifies feature sharing in GNNs via class‑conditional centroids and linear‑probe directions. Across architectures, width induces a three‑phase trajectory in overlap; topology shapes node‑level geometry that pooling re‑mixes toward the task; and sharper pooling drives axis alignment and reduces lever sharing. We also observed metastable low‑rank graph embeddings in shallow models. Practically, modest increases in width, LeakyReLU in the final MLP, and sharper but stable pooling improve interpretability without sacrificing accuracy. A key next step is to link superposition dynamics to over‑smoothing and over‑squashing, integrating our geometric view with spectral, dynamical, and topological analyses.
Appendix A Datasets
A.1 The Pairwise dataset
Graph structure and node features.
The Pairwise dataset consists of graphs that are simple chains of fixed length $L=20$. We represent the chain as an undirected path on nodes $\{1,\dots,L\}$ with edges $(i,i+1)$ for $i=1,\dots,L-1$, encoded as two directed edges $(i,i+1)$ and $(i+1,i)$; in our experiments we do not add self-loops. Each node $i$ carries a $k$-dimensional feature vector $x_{i}\in\{0,1\}^{k}$, where $k$ is the number of categories (we denote the corresponding dataset by Pairwisek). Node features are sampled independently as follows: first draw an activation flag $b_{i}\sim\mathrm{Bernoulli}(p)$ with $p=0.9$; if $b_{i}=1$, sample a category $c_{i}\sim\mathrm{Unif}\{1,\dots,k\}$ and set $(x_{i})_{c_{i}}=1$ and all other entries to 0, while if $b_{i}=0$ we set $x_{i}=\mathbf{0}$. Thus each node is either all-zero or one-hot, with a high expected fraction $p$ of active nodes. All random choices are made with a fixed seed ($42$) for reproducibility.
Label definition, deduplication, and split.
Given a feature matrix $X\in\{0,1\}^{L\times k}$ on the chain, the multi-label target $y\in\{0,1\}^{k}$ is defined coordinate-wise by
$y_{c}=\mathbf{1}\Big\{\exists\,(i,j)\in E\text{ with }(x_{i})_{c}=(x_{j})_{c}=1\Big\},\qquad c=1,\dots,k,$
where $E$ is the set of chain edges (self-loops, if present, are ignored when computing $y$). In words, class $c$ is active if and only if there exists at least one edge whose two endpoints both carry the one-hot category $c$. To build the dataset, we repeatedly sample node features $X$ as above, compute $y$, and retain only unique graphs, where uniqueness is defined by the node-feature pattern: we hash $X$ (after casting to uint8) and reject duplicates until we have $n_{\mathrm{train}}+n_{\mathrm{test}}=3000$ distinct samples. We then randomly shuffle these 3000 graphs (using the same seed $42$) and split them into $n_{\mathrm{train}}=2000$ training and $n_{\mathrm{test}}=1000$ test graphs, corresponding to a test ratio of $1/3$.
A.2 The Conjunction dataset
Data generation.
We instantiate the Conjunction dataset by drawing $n_{\text{train}}=3000$ and $n_{\text{test}}=1000$ graphs with a fixed random seed of $42$. For each graph we first sample a latent category $c\in\{\text{none},A\text{-only},B\text{-only},\text{both}\}$ from a categorical prior $(0.25,0.25,0.25,0.25)$. Conditioned on $c$, we specify the required counts $r_{\ell}\in\{0,1\}$ of cycles of lengths $\ell\in\{3,4,5,6\}$ as follows. For $A$-only we set $(r_{3},r_{4},r_{5},r_{6})=(1,0,0,1)$; for $B$-only $(0,1,1,0)$; for both $(1,1,1,1)$. For the “none” category we sample uniformly from the four non-target pairs $(1,0,1,0)$, $(0,1,0,1)$, $(1,1,0,0)$, $(0,0,1,1)$, which guarantees that neither conjunction $C_{3}\wedge C_{6}$ nor $C_{4}\wedge C_{5}$ holds. Given these required counts, we then draw independent extra copies $e_{\ell}\sim\mathrm{Binomial}(8,0.5)$ for each length $\ell$ with $r_{\ell}=1$, and set the total motif counts to $n_{\ell}=r_{\ell}+e_{\ell}$; lengths with $r_{\ell}=0$ remain absent. This corresponds to a symmetric Bernoulli prior over up to eight additional copies per present motif. The final labels are then computed deterministically from the counts as
$y_{A}(G)=\mathbf{1}\{n_{3}\geq 1\wedge n_{6}\geq 1\},\qquad y_{B}(G)=\mathbf{1}\{n_{4}\geq 1\wedge n_{5}\geq 1\},$
so that the four categories correspond to $(y_{A},y_{B})\in\{(0,0),(1,0),(0,1),(1,1)\}$ as intended.
Graph construction.
Given the counts $(n_{3},n_{4},n_{5},n_{6})$, we materialize a concrete graph by placing all motif instances on disjoint node sets. For each of the $n_{\ell}$ copies of a cycle of length $\ell$, we create a simple cycle on $\ell$ new nodes and designate one “anchor” node on that cycle. To increase local structural variety while preserving the core motifs, we attach “whisker” paths to each anchor: for every anchor we sample a number of whiskers $W\sim\mathrm{Unif}\{0,1,2\}$, and for each whisker we sample a length $L_{\text{wh}}\sim\mathrm{Unif}\{1,2\}$ and attach a simple path of length $L_{\text{wh}}$ starting from the anchor. Letting $\mathcal{A}$ denote the set of anchors of all motif instances, we then randomly permute $\mathcal{A}$ and connect consecutive anchors by simple paths of length $L_{\text{br}}\sim\mathrm{Unif}\{1,2\}$, ensuring that the overall graph is connected while the cycles themselves remain node-disjoint. Node features are one-dimensional and equal to the node degree (i.e., we use purely structural degree features with no learned or random attributes).
Deduplication and split.
To avoid distributional artefacts from re-using the same graph topology, we perform bucketed rejection sampling with Weisfeiler–Lehman (WL) deduplication. We maintain four buckets (one per category) and generate candidate graphs as above until each bucket contains the desired number of graphs implied by the class prior (subject to rounding), rejecting any candidate whose 1-WL hash (obtained from three iterations of 1-WL colour refinement on the undirected graph and hashing the resulting colour histogram) has already been seen in that bucket, with a cap of 60,00060{,}000 proposals per bucket. Once a total of $4000$ unique graphs have been collected, we perform a per-bucket random split into $3000$ training and $1000$ test graphs using the same seed ($42$), which preserves the class proportions up to integer rounding.
Appendix B Optimizer and training
B.1 Optimizer details
All models are trained with the Adam optimizer as implemented in PyTorch, using the default parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$ and $\epsilon=10^{-8}$. Unless otherwise stated, we use a mini-batch size of $256$, no learning-rate scheduling, and no early stopping; every configuration is run over multiple random seeds, but the optimizer hyperparameters are identical across seeds. No dropout is performed.
Table 2: Training hyperparameters (optimizer, learning rate, weight decay, and batch size) by experiment and architecture.
Experiment Task / dataset Model Learning rate Weight decay Epochs
1 Pairwise16 GCN 0.10 0 300
1 Pairwise16 GIN 0.01 0 150
1 Pairwise16 GAT 0.01 0 150
2 Conjunction GCN 0.001 $10^{-5}$ 1600
2 Conjunction GIN 0.001 $10^{-5}$ 1600
2 Conjunction GAT 0.001 $10^{-5}$ 1600
3 Pairwise16 GCN 0.10 0 200
3 Pairwise16 GIN 0.01 0 200
3 Pairwise16 GAT 0.01 0 200
4 Pairwise16 GCN 0.10 0 250
4 Pairwise16 GIN 0.01 0 250
4 Pairwise16 GAT 0.01 0 250
Experiment 1 (Dimension induced bottleneck on Pairwise).
We use mean global pooling, batch size $256$, and the learning rates and weight decays from Table 2. We train GCN models for $300$ epochs and GIN/GAT models for $150$ epochs:
$\text{GCN: }\text{lr}=0.10,~\text{wd}=0,\;300~\text{epochs};\qquad\text{GIN/GAT: }\text{lr}=0.01,~\text{wd}=0,\;150~\text{epochs}.$
The optimizer is Adam with the above defaults and no scheduler.
Experiment 2 (Topology induced superposition).
For the second experiment we again use Adam, batch size $256$, and mean pooling, but employ smaller learning rates and a small amount of weight decay. All architectures share the same training schedule:
$\text{GCN/GIN/GAT: }\text{lr}=10^{-3},\;\text{wd}=10^{-5},\;1600~\text{epochs}.$
Training is performed by calling model.fit(train_loader, optimizer, num_epochs=1600) with no additional regularization beyond weight decay.
Experiment 3 (alignment on Pairwise16).
In the alignment experiments on the Pairwise16 dataset we sweep the hidden dimension over $[10,16,22]$ with batch size $256$, and the same Adam learning rates and weight decays as in Experiment 1 (see Table 2). We sweep the generalized-mean pooling exponent over
$p\in\{1.0,2.0,4.0,8.0,p_{\max}\},$
where $p_{\max}$ denotes the largest exponent used in our code (approximating max pooling); all other optimization hyperparameters are kept fixed across $p$.
Experiment 4 (rank / selectivity on Pairwise16).
We use the ExperimentRank wrapper with hidden dimension $16$, batch size $256$, and the Adam settings from Table 2. Each model is trained via model.fit(train_loader, optimizer, num_epochs=num_epochs), where num_epochs is a fixed constant within this experiment and does not depend on the random seed; no learning-rate schedule or additional regularization is applied beyond the specified weight decay.
Appendix C Generalized Mean Pooling